# Streaming Generated Gaussian Process Experts for Online Learning and Control

### v2

@@ 1 Introduction
Real-time learning has become essential for modeling the dynamics of physical systems, where machine learning models must be continuously updated to adapt to changing environments while satisfying the requirements of responsiveness and safety. This capability is especially critical for autonomous systems, such as underwater vehicles, aerial drones, and healthcare robots, that operate in complex and safety-critical environments (Yang et al. 2025). In such scenarios, the ability to learn accurate models online and seamlessly integrate them into control loops is key to ensuring robust and efficient operation (Dai et al. 2023).
@@
Gaussian Processes (GPs) provide a powerful nonparametric framework for modeling dynamical systems, particularly in safety-critical applications, due to their ability to quantify uncertainty (Wachi et al. 2018). However, despite their modeling flexibility, standard GP inference suffers from severe scalability limitations. It scales cubically with the number of data points, requiring $\mathcal{O}(N^{3})$ time and $\mathcal{O}(N^{2})$ memory for $N$ training data. This scalability bottleneck makes conventional GPs impractical for online applications involving continuously streaming data or long-term deployments.
@@
To mitigate these limitations, a variety of scalable approximation techniques have been developed. Among global approximations, sparse Gaussian Processes introduce a set of $M\ll N$ inducing points to efficiently summarize the training data, reducing the complexity to $\mathcal{O}(NM^{2})$ for training and $\mathcal{O}(M^{2})$ for prediction. Notable methods include fully independent training conditional approximation (Snelson and Ghahramani 2005), variational free energy (Naish-guzman and Holden 2007), and latent projection techniques (Reeb et al. 2018). While streaming variants such as incremental sparse spectrum GP (ISSGP) (Gijsberts and Metta 2013) and streaming sparse GP (SSGP) (Bui, Nguyen, and Turner 2017) allow for incremental online updates, these frequently require computationally expensive optimization at each update, which can undermine their practical use in latency-sensitive settings. Moreover, these methods typically forgo guarantees on prediction error bounds, sacrificing reliability in critical applications.
@@
Distributed Gaussian Processes (DGPs) provide complementary scalability by partitioning data across multiple processors or agents for parallel processing. Within this paradigm, mixture-of-experts (MoE) approaches (Tresp 2000b; Yuan and Neubauer 2008; Trapp et al. 2020) aggregate predictions from independently trained GP experts using fixed or learned weights. Extensions such as product-of-experts (PoE) (Cohen et al. 2020), generalized PoE (gPoE) (Cao and Fleet 2015), and correlated PoE with sparse GPs (Schürch et al. 2023) have been proposed to better utilize uncertainty in aggregation. The Bayesian committee machine (BCM) (Tresp 2000a) and its robust variant (rBCM) (Deisenroth and Ng 2015; Liu et al. 2018) explicitly integrate the GP prior to mitigate overconfidence. To address real-time learning problems within distributed frameworks, Lederer et al. (2021) proposed LoG-GP, which incrementally constructs a tree-structured ensemble of GP experts offering a scalable solution for DGPs. However, the reliance on partitioning along a single dimension limits effectiveness in high-dimensional spaces. The structure of DGPs enables parallel and distributed deployment, with each GP model hosted on a separate computational node or agent. This is widely employed in multi-agent systems, where agents can independently operate local GPs and collaborate through a communication network for cooperative learning (Yang et al. 2021, 2024a, 2024b; Lederer et al. 2023). Nevertheless, these methods overlook online learning requirements, particularly the need to efficiently incorporate new observations and adapt to nonstationary environments (Yuan and Zhu 2024; Dai et al. 2024, 2025). While Yang, Dai, and Hirche (2025) tackle asynchronous communication, their computation-aware random splitting strategy fails to exploit spatial or temporal correlations in the streaming data.
@@ 2 Problem Statement and Preliminaries
Gaussian Processes represent a Bayesian non-parametric approach commonly employed for function regression and approximation tasks. It assumes that the target function $f(\cdot)$ is sampled from a GP prior, denoted as $f(\cdot)\sim\mathcal{GP}(m(\cdot),\kappa(\cdot,\cdot))$, where $m(\cdot)$ is the mean function (often set to zero when no prior knowledge is assumed) (Rasmussen and Williams 2006) and $\kappa(\cdot,\cdot)$ is the covariance kernel (e.g., the squared exponential kernel). With the streaming observations accumulated up to time step $t_{k}\in\mathbb{R}_{\geq 0}$ forming the dataset $\mathbb{D}(t_{k})=\{(\bm{x}^{s},y^{s})\}_{s=1,2,\dots,N(t_{k})}$, where $N(t_{k})=|\mathbb{D}(t_{k})|$, the posterior mean and variance functions of the GP can be derived as follows
@@
with the training output $\bm{y}=[y^{1},\dots,y^{N(t_{k})}]^{\top}$. The kernel matrix $\bm{K}$ is defined by its elements $K_{ij}=\kappa(\bm{x}^{i},\bm{x}^{j})$ (Rasmussen and Williams 2006).
@@ 3.1 Kernel-Induced Distance
To enable partitioning and localization for newly generated GP experts, we adopt a center-based representation of the training inputs, where the center of the $i$-th expert $\mathcal{GP}_{i}$ is defined as $\bm{c}_{i}=\frac{1}{N_{i}}\sum_{k=1}^{N_{i}}\bm{x}_{i}^{k}$, following (Nguyen-tuong, Peters, and Seeger 2008). Each expert maintains its own representative center to support fast and scalable online allocation. To facilitate online computation, the center of each expert is updated incrementally as new data points are assigned. Specifically, when a new point $\bm{x}^{k}$ arriving at time step $t_{k}$ is allocated to the selected expert, the expert center is updated as
@@ 3.2 Progressive Expert Generation Strategy
In conventional online GP approximation methods, new data points are typically assigned to the first available expert that has not yet reached its capacity (Nguyen-tuong, Peters, and Seeger 2008; Lederer et al. 2021; Yang, Dai, and Hirche 2025). Once the expert becomes full, it is either no longer updated or directly split into sub-experts, often without explicitly accounting for the underlying distributional properties and space features of the data. In contrast, our SkyGP framework dynamically allocates each incoming data point to the most appropriate expert based on kernel proximity and time-aware factor. Specifically, from the prediction phase, we have already calculated the kernel distances between the new arriving point and the experts within an adaptive window $W\in\mathbb{I}_{[1,\bar{W}]}$ and found the nearest experts with in the expert set $\mathcal{N}(t_{k})$, where $\mathcal{N}(t_{0})=1$ (see Algorithm 1). After locating the previous index $\vartheta_{\text{perv}}$ of the nearest expert, we are able to process the streaming data based on the index list $\mathcal{I}$ filtered by the time-aware factor $\vartheta\in(0,1]$ initialized by $1$, which encodes the usage history of the experts. If the expert has not reached its capacity, the data point is added to it (see Figure 1). Otherwise, the two SkyGP variants handle this situation in two different ways outlined in Algorithm 2.
@@ 4.1 Learning Performance of SkyGP
Compared with the other probabilistic (Srinivas et al. 2012; Scharnhorst et al. 2022) or deterministic bounds (Maddalena, Scharnhorst, and Jones 2021; Hashimoto et al. 2022), the error bound chosen in this paper follows (Lederer, Umlauft, and Hirche 2019), which results in an amplitude constant $\beta$ of posterior variance $\sigma(\cdot)$. Moreover, it is shown that there always exists a sufficiently small grid factor $\tau\in\mathbb{R}_{+}$ such that $\gamma(\bm{x})\ll\beta\sigma(\bm{x})$ for all $\bm{x}\in\mathbb{X}$, i.e., the posterior variance dominates the error bound $\eta(\bm{x})$. Therefore, the prediction error bound could also be written as $\eta(\bm{x})\leq 2\beta\sigma(\bm{x})$ (Lederer et al. 2024).
@@ A General Control System
This assumption indicates the asymptotic output stability of the equivalent system $\dot{\bm{x}}=\bm{\mu}(\bm{x},\bm{u})+\bm{g}(\bm{x},\bm{u})$ using the control law $\bm{\pi}(t,\bm{x},\bm{\mu})$. Compared to the condition in conventional asymptotic stability (Khalil 2002), the additional term $\|\nabla_{\bm{x}}V(t,\bm{x})\|^{2}/(4\varepsilon)$ introduced in (17) is used following input-to-state control Lyapunov functions to partially compensate for the potential effect from unknown disturbances. Later, an example is provided to demonstrate how to design a model-based controller that satisfies this assumption. Notably, the true dynamical system is unknown $\bm{f}(\cdot)$ instead of $\bm{\mu}(\cdot)$, whose performance is shown as follows.
@@
where $\bm{B}=[\bm{0}_{d_{q}\times d_{q}},\bm{M}^{-T}(\bm{q})]^{T}$. The control gains $\bm{K}_{p}$ and $\bm{K}_{d}$ are chosen such that $\bm{A}=\begin{bmatrix}\bm{0}_{m_{q}\times m_{q}}&\bm{I}_{m_{q}}\\ \bm{K}_{p}&\bm{K}_{d}\end{bmatrix}$ is Hurwitz. The matrix $\bm{P}$ is the solution to the Lyapunov equation associated with the quadratic Lyapunov function $V(t,\bm{z})=\bm{z}^{\top}\bm{P}\bm{z}$, where $\bm{P}\succ 0$ satisfies $\bm{A}^{\top}\bm{P}+\bm{P}\bm{A}=-\bm{Q}$ with $\bm{Q}\succ 0$ being a positive definite matrix. The existence and uniqueness of the solution $\bm{P}$ is guaranteed due to Hurwitz $\bm{A}$ (Khalil 2002). Then, the control performance in terms of tracking error bound is shown as follows.
@@ 5 Experimental Evaluation
We conduct comprehensive experiments to evaluate the proposed SkyGP framework from both regression and control perspectives. In Section 5.1, we assess the regression performance on multiple real-world benchmark datasets. Then, a closed-loop control performance comparison is evaluated against SOTA approaches on a nonlinear dynamical system in Section 5.2. Finally, we present an ablation study analyzing different algorithmic design choices, such as the use of Euclidean distance, dropped center, and more.
@@ 5.1 Regression Performance Evaluation
We evaluate the proposed SEGP framework on four real-world benchmark datasets. The SARCOS dataset contains 44,484 samples with 21 input features. The PUMADYN32NM (PUMA) dataset consists of 7,168 samples with 32 input features. The KIN40K dataset includes 10,000 samples with 8-dimensional inputs. Finally, the ELECTRIC dataset records household electricity consumption, comprising over 2 million samples with 11-dimensional inputs. For simplicity, we use the first 20,000 samples [^2] For additional results, including regression and control performance as well as ablation study analysis, refer to Appendix B.. To assess the performance of SkyGP, we compare our two variants against several strong baselines: LoG-GP (Lederer et al. 2021) and Local GPs (Nguyen-tuong, Peters, and Seeger 2008) with each expert holding at most 50 data points, SSGP with 50 inducing points (Bui, Nguyen, and Turner 2017), and ISSGP with 200 random features (Gijsberts and Metta 2013). All methods use an ARD kernel with hyperparameters pre-trained with the first 1000 samples in each dataset. Each model is evaluated in a sequential setting, where predictions and updates are performed on-the-fly as each new data point arrives. We vary the maximal number of experts $\bar{\mathcal{N}}\in\{1,2,4\}$ in all benchmark. Each expert with SkyGP holds at most $\bar{N}=50$ data points and the maximal search window size $\bar{W}=40$ and uses the rBCM aggregation method. We use the average prediction and update times, as well as the standardized mean squared error (SMSE) and the mean standardized log loss (MSLL) in a sequential interpretation (Lederer et al. 2021). We denote SkyGP-Fast with a maximum of one expert as SkyGP-F-1, and SkyGP-Dense with one expert as SkyGP-D-1. Other variants follow the same naming convention.
@@
Table 1 reports the average SMSE and MSLL across three benchmark datasets, demonstrating that the proposed SkyGP variants consistently outperform baseline methods such as LoG-GP and LocalGPs in both predictive accuracy and uncertainty calibration. While ISSGP achieves the best SMSE and MSLL on the PUMA dataset, its prediction and update time is over 30 times slower, as detailed in Table 2, making it less suitable for real-time applications. Notably, SkyGP-D-4 achieves the best overall performance, with the lowest SMSE (0.017) and MSLL (-2.03) on the SARCOS dataset, and competitive results on PUMA and ELECTRIC. Table 2 summarizes the average prediction and update times across three datasets. SkyGP-F-1 achieves the lowest overall latency, with update times consistently at 0.04s and the fastest prediction on SARCOS (0.16s), making it well-suited for real-time applications. SkyGP-D-1 has the best prediction time on ELECTRIC (0.14s) while maintaining competitive update performance. In contrast, traditional baselines such as ISSGP and SSGP exhibit significantly higher computational costs, with ISSGP requiring up to 18s for prediction and 7s for update on SARCOS, rendering them impractical for online deployment. Notably, SSGP fails to produce results on the ELECTRIC dataset within 20s, so we omit it.

### v3

@@ 1 Introduction
Real-time learning has become essential for modeling the dynamics of physical systems, where machine learning models must be continuously updated to adapt to changing environments while satisfying the requirements of responsiveness and safety. This capability is especially critical for autonomous systems, such as underwater vehicles, aerial drones, and healthcare robots, that operate in complex and safety-critical environments (yang_EAAI2025_safe). In such scenarios, the ability to learn accurate models online and seamlessly integrate them into control loops is key to ensuring robust and efficient operation (dai_L4DC2023_can).
@@
Gaussian Processes (GPs) provide a powerful nonparametric framework for modeling dynamical systems, particularly in safety-critical applications, due to their ability to quantify uncertainty (Wachi_2018AAAI_safe). However, despite their modeling flexibility, standard GP inference suffers from severe scalability limitations. It scales cubically with the number of data points, requiring $\mathcal{O}(N^{3})$ time and $\mathcal{O}(N^{2})$ memory for $N$ training data. This scalability bottleneck makes conventional GPs impractical for online applications involving continuously streaming data or long-term deployments.
@@
To mitigate these limitations, a variety of scalable approximation techniques have been developed. Among global approximations, sparse Gaussian Processes introduce a set of $M\ll N$ inducing points to efficiently summarize the training data, reducing the complexity to $\mathcal{O}(NM^{2})$ for training and $\mathcal{O}(M^{2})$ for prediction. Notable methods include fully independent training conditional approximation (Snelson_NeurIPS2005_Sparse), variational free energy (Naish_NeurIPS2007_Generalized), and latent projection techniques (Reeb_NeurIPS2018_Learning). While streaming variants such as incremental sparse spectrum GP (ISSGP) (Arjan_NN2013_Real) and streaming sparse GP (SSGP) (bui_NeurIPS2017_Streaming) allow for incremental online updates, these frequently require computationally expensive optimization at each update, which can undermine their practical use in latency-sensitive settings. Moreover, these methods typically forgo guarantees on prediction error bounds, sacrificing reliability in critical applications.
@@
Distributed Gaussian Processes (DGPs) provide complementary scalability by partitioning data across multiple processors or agents for parallel processing. Within this paradigm, mixture-of-experts (MoE) approaches (Tresp_NeurIPS2000_Mixtures; Yuan_NeurIPS2008_Variational; Trapp_AISTATS2020_Deep) aggregate predictions from independently trained GP experts using fixed or learned weights. Extensions such as product-of-experts (PoE) (Cohen_ICML2020_Healing), generalized PoE (gPoE) (cao2015generalizedproductexpertsautomatic), and correlated PoE with sparse GPs (Scheurch_ML2023_Correlated) have been proposed to better utilize uncertainty in aggregation. The Bayesian committee machine (BCM) (tresp_NC2000_bayesian) and its robust variant (rBCM) (Deisenroth_ICML2015_Distributed; liu_ICML2018_Generalized) explicitly integrate the GP prior to mitigate overconfidence. To address real-time learning problems within distributed frameworks, lederer_ICML2021_Gaussian proposed LoG-GP, which incrementally constructs a tree-structured ensemble of GP experts offering a scalable solution for DGPs. However, the reliance on partitioning along a single dimension limits effectiveness in high-dimensional spaces. The structure of DGPs enables parallel and distributed deployment, with each GP model hosted on a separate computational node or agent. This is widely employed in multi-agent systems, where agents can independently operate local GPs and collaborate through a communication network for cooperative learning (yang_CDC2021_Distributed; yang_AAMAS2024_whom; yang_ACC2024_cooperative; Lederer_TAC2024_cooperative). Nevertheless, these methods overlook online learning requirements, particularly the need to efficiently incorporate new observations and adapt to nonstationary environments (Yuan_TAC2024_Lightweight; dai_EJC2024_Decentralized; dai_TNNLS2025_Cooperative). While yang_AAAI2025_Asynchronous tackle asynchronous communication, their computation-aware random splitting strategy fails to exploit spatial or temporal correlations in the streaming data.
@@ 2 Problem Statement and Preliminaries
Gaussian Processes represent a Bayesian non-parametric approach commonly employed for function regression and approximation tasks. It assumes that the target function $f(\cdot)$ is sampled from a GP prior, denoted as $f(\cdot)\sim\mathcal{GP}(m(\cdot),\kappa(\cdot,\cdot))$, where $m(\cdot)$ is the mean function (often set to zero when no prior knowledge is assumed) (rasmussenGaussianProcessesMachine2006) and $\kappa(\cdot,\cdot)$ is the covariance kernel (e.g., the squared exponential kernel). With the streaming observations accumulated up to time step $t_{k}\in\mathbb{R}_{\geq 0}$ forming the dataset $\mathbb{D}(t_{k})=\{(\bm{x}^{s},y^{s})\}_{s=1,2,\dots,N(t_{k})}$, where $N(t_{k})=|\mathbb{D}(t_{k})|$, the posterior mean and variance functions of the GP can be derived as follows
@@
with the training output $\bm{y}=[y^{1},\dots,y^{N(t_{k})}]^{\top}$. The kernel matrix $\bm{K}$ is defined by its elements $K_{ij}=\kappa(\bm{x}^{i},\bm{x}^{j})$ (rasmussenGaussianProcessesMachine2006).
@@ 3.1 Kernel-Induced Distance
To enable partitioning and localization for newly generated GP experts, we adopt a center-based representation of the training inputs, where the center of the $i$-th expert $\mathcal{GP}_{i}$ is defined as $\bm{c}_{i}=\frac{1}{N_{i}}\sum_{k=1}^{N_{i}}\bm{x}_{i}^{k}$, following (Nguyen_NeurIPS2008_Local). Each expert maintains its own representative center to support fast and scalable online allocation. To facilitate online computation, the center of each expert is updated incrementally as new data points are assigned. Specifically, when a new point $\bm{x}^{k}$ arriving at time step $t_{k}$ is allocated to the selected expert, the expert center is updated as
@@ 3.2 Progressive Expert Generation Strategy
In conventional online GP approximation methods, new data points are typically assigned to the first available expert that has not yet reached its capacity (Nguyen_NeurIPS2008_Local; lederer_ICML2021_Gaussian; yang_AAAI2025_Asynchronous). Once the expert becomes full, it is either no longer updated or directly split into sub-experts, often without explicitly accounting for the underlying distributional properties and space features of the data. In contrast, our SkyGP framework dynamically allocates each incoming data point to the most appropriate expert based on kernel proximity and time-aware factor. Specifically, from the prediction phase, we have already calculated the kernel distances between the new arriving point and the experts within an adaptive window $W\in\mathbb{I}_{[1,\bar{W}]}$ and found the nearest experts with in the expert set $\mathcal{N}(t_{k})$, where $\mathcal{N}(t_{0})=1$ (see Algorithm 1). After locating the previous index $\nu_{\text{prev}}$ of the nearest expert denoted by $\nu_{\text{nr}}$, we are able to process the streaming data based on the index list $\mathcal{I}$ filtered by the time-aware factor $\vartheta\in(0,1]$ initialized by $1$, which encodes the usage history of the experts. Moreover, we define the upper bound of $\vartheta$ as $\bar{\vartheta}$ to filter out the outdated experts. If the expert has not reached its capacity, the data point is added to it (see Figure 1). Otherwise, the two SkyGP variants handle this situation in two different ways outlined in Algorithm 2.
@@ 4.1 Learning Performance of SkyGP
Compared with the other probabilistic (srinivas2012information; scharnhorst2022robust) or deterministic bounds (maddalena2021deterministic; hashimoto2022learning), the error bound chosen in this paper follows (lederer2019uniform), which results in an amplitude constant $\beta$ of posterior variance $\sigma(\cdot)$. Moreover, it is shown that there always exists a sufficiently small grid factor $\tau\in\mathbb{R}_{+}$ such that $\gamma(\bm{x})\ll\beta\sigma(\bm{x})$ for all $\bm{x}\in\mathbb{X}$, i.e., the posterior variance dominates the error bound $\eta(\bm{x})$. Therefore, the prediction error bound could also be written as $\eta(\bm{x})\leq 2\beta\sigma(\bm{x})$ (lederer2024safe).
@@ A General Control System
This assumption indicates the asymptotic output stability of the equivalent system $\dot{\bm{x}}=\bm{\mu}(\bm{x},\bm{u})+\bm{g}(\bm{x},\bm{u})$ using the control law $\bm{\pi}(t,\bm{x},\bm{\mu})$. Compared to the condition in conventional asymptotic stability (khalil2015nonlinear), the additional term $\|\nabla_{\bm{x}}V(t,\bm{x})\|^{2}/(4\varepsilon)$ introduced in (17) is used following input-to-state control Lyapunov functions to partially compensate for the potential effect from unknown disturbances. Later, an example is provided to demonstrate how to design a model-based controller that satisfies this assumption. Notably, the true dynamical system is unknown $\bm{f}(\cdot)$ instead of $\bm{\mu}(\cdot)$, whose performance is shown as follows.
@@
where $\bm{B}=[\bm{0}_{d_{q}\times d_{q}},\bm{M}^{-\top}(\bm{q})]^{\top}$. The control gains $\bm{K}_{p}$ and $\bm{K}_{d}$ are chosen such that $\bm{A}=\begin{bmatrix}\bm{0}_{m_{q}\times m_{q}}&\bm{I}_{m_{q}}\\ \bm{K}_{p}&\bm{K}_{d}\end{bmatrix}$ is Hurwitz. The matrix $\bm{P}$ is the solution to the Lyapunov equation associated with the quadratic Lyapunov function $V(t,\bm{z})=\bm{z}^{\top}\bm{P}\bm{z}$, where $\bm{P}\succ 0$ satisfies $\bm{A}^{\top}\bm{P}+\bm{P}\bm{A}=-\bm{Q}$ with $\bm{Q}\succ 0$ being a positive definite matrix. The existence and uniqueness of the solution $\bm{P}$ is guaranteed due to Hurwitz $\bm{A}$ (khalil2015nonlinear). Then, the control performance in terms of tracking error bound is shown as follows.
@@ 5 Experimental Evaluation
We conduct comprehensive experiments to evaluate the proposed SkyGP framework from both regression and control perspectives. In Section 5.1, we assess the regression performance on multiple real-world benchmark datasets. Then, a closed-loop control performance comparison is evaluated against SOTA approaches on a nonlinear dynamical system in Section 5.2. Finally, we present an ablation study analyzing different algorithmic design choices [^2] For additional results and ablation study analysis, refer to Appendix B..
@@ 5.1 Regression Performance Evaluation
We evaluate the proposed SkyGP framework on four real-world benchmark datasets. The SARCOS dataset contains 44,484 samples with 21 input features. The PUMADYN32NM (PUMA) dataset consists of 7,168 samples with 32 input features. The KIN40K dataset includes 10,000 samples with 8-dimensional inputs. The ELECTRIC dataset records household electricity consumption, comprising over 2 million samples with 11-dimensional inputs. For simplicity, we use the first 20,000 samples of ELECTRIC dataset. To assess the performance of SkyGP, we compare our two variants against several strong baselines: LoG-GP (lederer_ICML2021_Gaussian) and Local GPs (Nguyen_NeurIPS2008_Local) with each expert holding at most 50 data points, SSGP with 50 inducing points (bui_NeurIPS2017_Streaming), and ISSGP with 200 random features (Arjan_NN2013_Real). All methods use an ARD kernel with hyperparameters pre-trained with the first 1000 samples in each dataset. Each model is evaluated in a sequential setting, where predictions and updates are performed on-the-fly as each new data point arrives. We evaluated the average prediction and update times, as well as the standardized mean squared error (SMSE) and the mean standardized log loss (MSLL) in a sequential interpretation (lederer_ICML2021_Gaussian). The maximal number of experts $\bar{\mathcal{N}}\in\{1,2,4\}$ are used in all benchmark. Each expert with SkyGP holds at most $\bar{N}=50$ data points and the maximal search window size $\bar{W}=40$ and uses the rBCM aggregation method. Set the time-aware decay weighting factor $\rho=0.995$ and $\bar{\vartheta}=10^{-3}$. SkyGP-Fast with a maximum of one expert is denoted by SkyGP-F-1, and SkyGP-Dense with one expert by SkyGP-D-1. Other variants follow the same naming convention.
@@
Table 1 shows the average SMSE and MSLL across three benchmark datasets, demonstrating that the proposed SkyGP variants consistently outperform baseline methods such as LoG-GP and LocalGPs in both predictive accuracy and uncertainty calibration. While ISSGP achieves the best SMSE and MSLL on the PUMA dataset, its prediction and update time is over 30 times slower, as detailed in Table 2, making it less suitable for real-time applications. Notably, SkyGP-D-4 achieves the best overall performance, with the lowest SMSE (0.017) and MSLL (-2.03) on the SARCOS dataset, and competitive results on PUMA and ELECTRIC. Table 2 summarizes the average prediction and update times across three datasets. SkyGP-F-1 achieves the lowest overall latency, with update times consistently at 0.04s and the fastest prediction on SARCOS (0.16s), making it well-suited for real-time applications. SkyGP-D-1 has the best prediction time on ELECTRIC (0.14s) while maintaining competitive update performance. In contrast, traditional baselines such as ISSGP and SSGP exhibit significantly higher computational costs, with ISSGP requiring up to 18s for prediction and 7s for update on SARCOS, rendering them impractical for online deployment. Notably, SSGP fails to produce results on the ELECTRIC dataset within 20s, so we omit it.